A sphere of mass m is ropped from a height h above the ground. on reaching theground it penetrates through a distance S and finally stops. the average resistance offered bythe sand to oppose the motion is
 

Grade:11

1 Answers

Susmita
5 years ago
At height h it has a initial speed u=0.Just before hitting the ground its speed v is
v2=u2+2fs=2gh
or,v=root(2gh)
So it hits the ground with an initial speed u=root(2gh) and after covering a distance s it comes to rest,ie,its final velocity v=0.If f is its acceleration inside ground then
v2=u2+2fs
Or,0=2gh+2fs
Or,f=-gh/s
So the resistance(force) offered by sand is
mf=mgh/s
Hope this helps